Astronomy Report – March 4th

Crew 223 Astronomy Report Mar 04 2020

Name: Florian Delpech
Crew: 223
Date: 3/4/2020

Robotic Telescope Requested (choose one) MDRS-14

Objects to be Imaged this Evening: Same as yesterday : IC 342 ; NGC 2775; NGC 4449 ; NGC 4254 ; NGC 4125 ; NGC 3338 ; NGC 2683 ; M100 ; NGC 6503; NGC 5248 ; NGC 4038. This is part of the research project aiming at observing about 30 galaxies per week to try to find a supernovae.

Images submitted with this report: A picture of M81.

Problems Encountered: When I wanted to calibrate images with AstroImageJ for M81 and IC1805, I used calibration files (MDRS-WF_mbias, MDRS-WF_mdark, etc.) taken from Skynet on March, 3rd. However, when I launched CCD Data Processor, the following error message appeared : "Fileinfo parameter error. Offset + image size > file length".

To solve the problem, I took calibration files from a previous observation (a week ago) and it worked. So maybe the calibration files from yesterday were too heavy.

Also, I don’t know if it is linked but the image with Halpha filter for both objects was very noised and so unusable.


Solar Features Observed: N/A

Images submitted with this report: N/A

Problems Encountered: N/A

Journalist Report – March 4th

Crew 223 Journalist Report 04Mar2020
Author: Clément Plagne, Crew Journalist
Sol 3
Title: The Commander’s Day

Today, our commander was busy. He was not lucky enough to go on EVA today, being instead stuck inside the Hab on HabCom duty. He’s helped out everywhere, from assisting the harvest in the GreenHab to lending a hand in the RAM to troubleshoot the recent spacesuit issues. Also, he’s the one tasked with doing the dishes after dinner tonight. It’s also his birthday. For the second year in a row now, Aurélien got a year older inside the station. Being away from family, home, and, to be quite fair, the whole world on a day like this is never easy, especially with a full
day’s worth of work to do. We’re great company, there’s no doubting it, but I don’t think any one of us is pretentious enough to think that’s sufficient. Besides, for what it’s worth, he’s been telling us he’s beginning to forget what birthdays outside the MDRS are like. Still, we do our best to make this a great day for him, starting with our EVA, that went very smoothly for everyone involved. For once, no spacesuit trouble to report, besides the fact that we had to use the slightly less comfortable ones, and were glad to lay down by the end of it.

The next step to a good day, as always for us, is a lovely meal. Our birthday boy himself, 23 years young now, took part in the making of some fantastic pizzas, tested and approved by all. Marion baked an absolutely wonderful banana bread – in what was possibly the toughest moment of the mission so far, we took the decision of waiting a bit before eating the leftovers. Sometimes I feel like all I’m talking about is food. When things get unusual or tough, you get back to the fundamentals: sleeping and eating. Both are going well for all, a great sign that the rest also works.

We actually ended up having some time before getting back to work to sit down at the table and play a little game of bluff and strategy. The dynamics of the team really is beginning to show, and we had some great fun playing (no one must tell Aurélien that we let him win). For the
third day running, Marion, our Crew Scientist, has had us run an experiment on communication in foreign languages. One person has the instruction for a LEGO build, and the other must construct the object from the pieces without either one seeing the other, with limited time. All teams are making great progress, going from Frankensteinesque monstrosities to things that actually begin to resemble the actual object! To add to that, our Health and Safety Officer Blandine has also continued with team building and relaxation exercises, and it’s once again been a lovely moment together and a great time to talk about how life is going so far in the station.

We’ve got a couple other things planned for Aurélien later on tonight, but we can’t spoil the surprises, so it’ll have to stay secret!

GreenHab Report – March 4th

Crew 223 GreenHab Report – 04-03-2020
GreenHab Officer: Valentin BELLEMIN-LAPONNAZ

Environmental control: Ambient with Door open

Average temperature: 23.6°C
Low temperature: 20°C
High temperature: 26°C

Hours of supplemental light: 4 hours (from 9 p.m to 1 a.m.)

Daily water usage for crops: 21.2 gallons

Daily water usage for research and/or other purposes: 1 gallon

Water in Blue Tank— 180.2 gallons

Time(s) of watering for crops: 4 (2 only for tomatoes)

Changes to crops: I transplanted the cucumber in -to 4 separated pots. I also transplanted the radish into 6 different pots for the experiment Music For Plants.

Narrative: This morning, Marion LEBRUN took care of the GreenHab, as I were on an EVA. She watered all the plants as usual. For lunch, I harvested Sage, Oregano and Basil to make some pizzas. This afternoon, with Blandine GORCE, we transplanted the fours plants of cucumbers which were together into 4 separated pots. We notices that two of them have fruits. We hope that we’ll be able to eat fresh cucumbers after the end of our mission. The experiment Music For Plants dried in the GreenHab. I had to find another solution. I decided to carry out the experiment with Radish. I transplanted some from the big blue tank into sported pots. I
hope they’ll survive from the change.

Harvest: 2g Basil / 2g Sage / 4g Oregano

Support/supplies needed: NA

Sol Summary – March 4th

Crew 223 Sol Summary Report 04-03-2020

Sol: 3 Summary Title: I’m getting old

Author’s name: Aurélien Mure – Crew Commander

Mission Status: Crew 223 newcomers are taking responsibilities

Sol Activity Summary:

The crew woke up at 6:50am and every crewmember wished me a happy birthday, which is really pleasant to start the day. The sport session began at 7:00am. The crew physical status are getting better and everybody is now used to the workout session. The first EVA without me
has been prepared really well by all the crew. They were really autonomous and cautious. They walked to Hab Ridge. They suffered from back pain but nothing unbearable.

For the lunch I prepared for my birthday meal two pizza, and Marion cooked banana bread. With this crew, no fear of bad cook!

The crew took 30 min to play a little strategy game after lunch. Luc and I fixed the Spacesuit 10 and found solutions for SpaceSuit 7. Florian worked on the solar oven and Valentin in the GreenHab. We went through the daily session of our Human Factor experiment and the
relaxation session.

The team spirit is really outstanding in here.

Look Ahead Plan:

The long range EVA planned tomorrow is set to go further into exploration. The water monitoring system struggles to work so we need to spend time on it.

Anomalies in work: None.

Weather: Sunny and warm all day (-6°C at 7:00am, about 8°C at 11:00am, 10°C at 6:00pm) 0m/s of wind all day

Crew Physical Status: All crew members safe and sound.

EVA: HabRidge exploration completed

Reports to be filed: Journalist Report, Operations report, EVA Request,
Green Hab report, HSO Report, Astronomy report, Science report

Support Requested: None

EVA Report – March 4th

Crew 223 EVA Report 04-03-2020
EVA # 3
Author: Blandine Gorce (XO)
Purpose of EVA: Change batteries of LOAC and MegaAres, and get a good view of our planet Mars from Hab Ridge

Start time: 9:03 AM
End time: 10:34 AM

Narrative: We walked out of the Hab at 09:08 AM after the five minutes of pressurization. We took the rovers to go and change the batteries of Mega Ares and LOAC. We were really organized and at 09:28 we had finished the procedures for the two measure instruments, so we came back to park the rovers next to the Hab.

We then started the exploration towards Hab Ridge, we walked a little bit next to the Hab and then we started to go up the ridge. It was not too hard because the soil was very dry and it was not too hot yet. When we arrived the top, the view was beautiful and we were really happy
to see the campus from there. We then decided to use Hab ridge road to go down the hill. we took the road close to North Ridge so we could get a good view at North Ridge which was beautiful.

We decided to come back early because the backpack were hurting a bit the back of one of us. We came back at the Hab at 10:29 AM to start the 5 minutes of pressurization.

Destination: The flat place next to the Hab and Hab Ridge

Coordinates (use UTM NAD27 CONUS):12S 518500 4250800, 12S 517850 4250350

Participants: Team red: Blandine Gorce (EVA Leader, HSO), Clément Plagne(Journalist), Luc Fortin(Engineer), Valentin Bellemin-Laponnaz(Green Hab Officer)

Road(s) and routes per MDRS Map: Entrance Road, Cow Dung Road

Mode of travel: Rovers : Curiosity & Deimos

Operations Report – March 4th

Crew 223 Operations Report 04/03/2020

SOL: 3

Name of person filing report: Luc FORTIN

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Tuesday generator:
Off at 5:00am
On at 9:10pm SOC was 84%

Wednesday generator:
Off at 6:00am
On at 9:00pm SOC was XX%

Diesel Reading – 70%

Station Propane Reading – 80%

Ethanol Free Gasoline – 7 gallons

Water (loft tank): 47 gallons

Water Meter: 149694.75 units

Water (static tank) – 334 gallons

Static to Loft Pump used – yes

Water in GreenHab – 180.2 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: yes
Hours: 149.5
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Spirit rover used: no
Hours: 133.6
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes.
Hours: Nothing to report
Beginning charge: Nothing to report
Ending charge: Nothing to report
Currently charging: Nothing to report

Curiosity rover used: yes
Hours: 139.3
Beginning charge: 100%
Ending charge: 91%
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (Honda, 1, 2 and 3)
Reason for use: Run during the engineer EVA for battery
Oil Added? no
ATV Fuel Used: Nothing to report
Hours the ATVs were Used today: 5 minutes
Notes on ATVs: Nothing to report

HabCar used and why, where? Nothing to report

CrewCar used and why, where? The CrewCar is in Grand Junction.

General notes and comments: Nothing to report

Summary of internet: All ok.

Summary of suits and radios: Suits 2, 3, 4, 9 were used during the EVA.
Suit 10 is working, and we will use it tomorrow.

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: We continue to fix suit 7 but haven’t
finished now. We first found a loose electrical connection at the
connection with the battery at the minus pod. We corrected it. But there
is still a problem with the suit, and we tend for another loose
electrical connection in the control box.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Science Report – March 4th

Research report 4 March 2020
Crew 223
Submitted by Crew Scientist Marion Lebrun

1. LOAC and Weather Station:
Everything is working as it should. SOL 0 was very windy and both experiments remained standing. Therefore, with the strong gusts of wind, the collected data should be interesting to analyse.

2. MegaARES:
Everything is working as it should. We have collected a first round of data; we have now 2.4Gb of Earth electric field data which is very satisfying.

3. Solar oven:
The oven is built (still in the Hab) and last night Florian downloaded the programs needed for the sensors. The sensors need to be tested before setting the oven on Mars.

4. BackPack and LCVG:
Nothing to report.

5. Foreign languages (Human Factors):
3 pairs have been formed : Aurélien and Florian speaking Spanish, Blandine and Clément speaking English, Valentin and Luc speaking German. Although speaking in a foreign language might have been disturbing at first for some of the crew members, after 2 days of experiment (SOL 1 and SOL 2), progress is definitely noticeable. Communication is clearly improving between the two members of each team; it was more efficient on SOL 2 day and the teams managed to build assemble more parts of LEGO than they did on SOL 1 – which is very encouraging.

6. Group collaboration (Human Factors):
Nothing to report.

7. Crew members’ individual planning display:
Blandine and Marion have decided what needs to be displayed: a dynamic schedule of the current sol, a window with Mission Support email refreshed every hour, the sport schedule for every member of the crew, and the available Internet data remaining. Blandine and Marion are
finishing their training on html before starting to actually code the app.

8. Relaxation and positive psychology (Human Factor):
Relaxation seems to have a greater impact on our crew than it had on Crew 222. As for positive psychology, the technique seems to be strengthening our crew’s cohesion.

9. Music for plants (GreenHab):
It is very hard to grow seeds in vitro since the atmosphere is too dry. This is why Crew 222 GreenHab Officer failed to grow her seeds in gelatin and why Valentin also failed to grow his on humidified paper towel. Valentin has thus decided to test his musical protocol on existing radish of the GreenHab growing in the ground.

10. Spirulina growth (GreenHab):
Everything is working as it should. We have already eaten a few grams of spirulina and made a banana and spirulina smoothie. We are now studying the impact of the type of fertilizer on the spirulina growth.

11. Aerospring (GreenHab):
Nothing to report.

12. Water consumption monitoring:
Nothing to report.

Copyright © The Mars Society. All rights reserved. | Main Site